Prev Next

SW&IT Meeting Minutes 2019-02-04

Attendees: Andrew, David, Matthieu, Petr, Stephen, Yang

Outlook

Labor Reporting

Report auto-generated every week, Do not edit this table, use entries in smartsheets.

20190114
LR20190114 - Bec, Matthieu 72.0
LR20190114 - Colson, Andrew 80.0
LR20190114 - Edmeades, David 56.0
LR20190114 - Sola, Igor 72.0
LR20190114 - Kubanek, Petr 73.0
LR20190114 - Hooper, Stephen 72.0
LR20190114 - Zhang, Yang 72.0
LR20190114 - Zhang, Xianyu 80.0
20190128
LR20190128 - Bec, Matthieu 40.0
LR20190128 - Colson, Andrew 40.0
LR20190128 - Edmeades, David 0.0
LR20190128 - Sola, Igor 30.5
LR20190128 - Kubanek, Petr 42.0
LR20190128 - Hooper, Stephen 80.0
LR20190128 - Zhang, Yang 40.0
LR20190128 - Zhang, Xianyu 40.0

News from management

Previous Action Items

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • allsky images named with timestamp
  • BCU f/w upgrade documentation
  • Dust monitor - bring back
  • AO improvement: Luci kwds
  • AO cabinet switch upgrade - follow up
  • wifi summit upgrade
  • Tucson CR power overhaul plan
  • n507/obs upgrade
  • abell / jhill ws
  • 2019 conferences - who where?

Team reports

DE

  • Safety training completed
  • More safety training completed
  • IT safetey training completed
  • VM Host network redundancy tests done; both 1 and 10Gb networks are redundant. Should be good to relocate them and begin work on Rack 3.
  • 6x 4TB disks wiped/checked for bad blocks and given to Stephen for new remote room computers. Continuing to check remaining drives.
  • Perle switch techopsy conclusion: blowed up real good. Ordered replacement. Do we know what power they applied and how? Note: units do not specify acceptable Voltage on the chassis, but the spec is 9.6 to 60 VDC.
  • Peter asked for help with an external drive from the mountain containing pictures from John Little. Unfortunately the disk has a physical problem and is inaccessable.
  • Ketiv/Vault permissions
Next:
  • VM Windows for Andrew and John Morris
  • Server room power document

PK

  • more C++11 (emplace, ..)
  • GCS changes - external scripts for image generation, GuidingThread reorganization, C++11 Sextractor interface (from ~500 lines to ~100)
  • collimation&pointing dissussions with Peter Grey
next:

AC

  • Last week
    • OAC
      • Added updated EZCA extension
    • Safety/Security training

  • This week
    • Delta-tau UMAC training
    • Particle monitor
      • contact LASAIR about firmware, etc.
      • re-build software
    • OAC
      • rebuild with EZCA for upcoming Eng night
    • Safety/Security training

MB

  • OC, WPC, management
  • automate OAC conf to git
  • looking at new webcomponents
  • shipped AHS to Shark-NIR
  • WIP with github issues and projects
  • ovms-tt meetings with Martin Gluck in Tucson
  • lbti pre/post meeting
  • safety trainings
next
  • delta-tau training
  • security training
  • setup my TCS simulator

YZ

  • Started LBTO online training for ComplianceAtWork (but haven't finished yet).
  • Finished UA's security awareness training.
  • IT#7285 MODS-1 Data Computer Froze mid-exposures, no response yet from Rick Pogge (hopefully their other MODS upgrades plan will address this problem).
  • AO-Cravat:
    • Tweaked several interface design (JS9 windows, keyboard bindings etc.), but more needs to be done (ongoing).
    • Created a fake MongoDB to test database design and more complex query (ongoing work).
    • Decided to use Google Chart as a chart plotting library since it integrates well with Google Polymer.
  • New TCS work:
    • Goal is to add new commands in IIF and new instrument authorization.
    • Currently working on setting up a new TCS dev environment and getting into the TCS parts I needed to know.

XZ

  • write the questions to Arcetri and Mircrogate for the power backplane
  • change the FLAO/SOUL SW for OVMS DSP
  • write the OVMS SW for configuring SwitchBCU and save OVMS telemetry
  • config the OVMS SW
  • help Alessandro packing lab set to the mountain
  • thanks to Stephen installed Matlab on the AdSec machines
  • burn the firmware in Tucson and mountain lab, and on the telescope
  • ovms-tt meetings
  • AdSec power off unexpected issue
  • prepare SwitchBCU replcement document
  • M2 hexpod limit IT
next
  • ovms
  • power backplane documents
  • prepare the Italy trip

IS

Last week:
  • LUCI
    • Added new subwindow coordinate to database and take a backup from the AFC coordinates table.
  • OT
    • ESM/ARGOS/SL/DL conversion script.
      • Michelle like it and discussed a posible more binocular generalization.
    • Discussion of LBTI PIT and starting developement.

SH

  • requested move of VxWorks crate into engineering lab (space) -- engineering is cool with this. Talk about buying a 19-inch rack to hold some equipment.
  • OVMS-TT host on mountain -- network issues, computer issues, hardware issues, BIOS issues, power issues.
  • Supported LBTI SOUL plumbing.
  • RPM for CA (done), RPM for wildcard (done).
  • Bucki laptop for "October" update to fix Windows Update (scary, scary!)
  • Matlab onto some machines
  • David C's safety fun. I now am certified not to electrocute myself.
  • update OAC git.

next:
  • WiFi on mountain
  • re0using old Synology drives in new rm507 machines.
  • self-service certificates
  • qemu-kvm under Fedora Atomic -- engineering server.
  • laptops/VMs for Delta-Tau? Need some info.
  • I should upgrade Matlab -- but I just don't feel it.
  • RPM for github swg private key and setup.
Topic revision: r20 - 04 Feb 2019, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback